OpenStudy (anonymous):

Rose worked the following problem incorrectly. Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer? Line 1) 30 ÷ 5(3 − 1) − 4 Line 2) 30 ÷ 5(2) − 4 Line 3) 30 ÷ 10 − 4 Line 4) 3 − 4 Line 5) −1 someone help?plz

OpenStudy (anonymous):

weird I feel like its correct..let me look at it closer -_-

OpenStudy (anonymous):

Ughh they seem all correct also.

OpenStudy (anonymous):

I'm only 70% sure, but I believe in this problem you are supposed to go from left to right after line 2 The problem at Line 2 looks like \[30\div 5 \times (2) - 4\]

OpenStudy (anonymous):

Ahh yeah the first line.

OpenStudy (anonymous):

BEDMAS for first line makes it the wrong one.

OpenStudy (anonymous):

ah I think Line 3 is where the mistake was made, it should become \[6 \times (2) -4\] From http://en.wikipedia.org/wiki/Order_of_operations in the Exceptions to the Standard section

OpenStudy (anonymous):

aight thnks:)

OpenStudy (anonymous):

It's line 1 !! x_x

OpenStudy (anonymous):

kinda confused line 1 or 3?

OpenStudy (anonymous):

Line 1 is the problem itself, you cant have a mistake there -_-

OpenStudy (anonymous):

30 ÷ 5(3 − 1) 5(3-1) = 15-5 30 ÷ 15-5 2-5= -3

OpenStudy (anonymous):

-4***

OpenStudy (anonymous):

= -2**

OpenStudy (anonymous):

2-4 = -2

OpenStudy (anonymous):

Let me retype it, I meesed up the whole thing.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Alright let me get some more opinions

OpenStudy (anonymous):

Mhmm... Unless brackets stay after.